# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v2

@@ IV-B2 Evaluation Metrics
For evaluating model performance across different CD subtasks, we employ a set of metrics tailored to the specific requirements of BCD, SCD, and BDA. In evaluating the BCD performance of the model, we use the following metrics:
@@
Cohen’s Kappa (KC)): Assessing the agreement between two raters beyond chance, Cohen’s Kappa is given by
@@
For BDA, we follow the protocol of xView2 Challenge [51] to split BDA into two subtasks, i.e., building localization and damage classification tasks. The following metrics are adopted to measure the performance of each subtask and the overall BDA performance:
F1 Score for Building Localization (): This metric assesses model’s accuracy in building localization task.
@@
Overall F1 Score (): This reflects the combined performance of the model in the tasks of building location and damage classification. It is defined as
@@ IV-C Comparison Methods
To show the superiority of our Mamba-based architecture in CD tasks, we select a number of of the state-of-the-art methods based on CNN and Transformer architectures for benchmarking across three distinct tasks: BCD, SCD, and BDA.
For BCD, our comparison methods includes a suite of CNN-based methods: FC-EF [22], FC-Siam-Diff [22], FC-Siam-Conc [22], SiamCRNN [17] (utilizing ResNet architectures ranging from ResNet-18 to ResNet-101 [67]), SNUNet [40], DSIFN [23], HANet [41], and CGNet [42]. For the Transformer-based methods, we compare against ChangeFormer (versions 1 to 6) [30], BIT (with ResNet-18 to ResNet-101 as the backbone for feature extraction) [28], TransUNetCD [33], and SwinSUNet [32].
@@ IV-D2 SCD
Based on the SCD results, the detailed “from-to” semantic change information can be obtained, which is presented in Fig. 10. Firstly, we can find that the “from-to” transition matrices obtained by our methods are very close to the ground truth, demonstrating the effectiveness of our MambaSCD architecture again. Then, based on the contents of matrices, we can find that the three semantic changes with the highest percentage between the two temporal phases are “non-vegetated ground surface building”, “low vegetation building”, “low vegetation non-vegetated ground surface”. These results indicate that the areas covered by the SECOND dataset was in the process of urbanization during this period.
@@
Table X further lists the number of parameters and GFLOPs of these methods. Compared to the Transformer-based ScanNet method, the GFLOPs of our three architectures are lower. Compared to SMNet, another Transformer-based method, the proposed MamabaSCD-Tiny can achieve higher SeK values with lower number of parameters and GFLOPs. These results demonstrate the effectiveness of the proposed methods and the potential of the Mamba architecture on the SCD task.
Fig. 11 shows some semantic change maps on the test set of the SECOND dataset. Due to land-cover category combinations, SCD often involves a considerable number of categories. For ease of visualization, we follow the visualization way in [24, 68, 20], namely only plotting the land-cover categories of the changed regions in both pre- and post-event images. In this way, the “from-to” semantic change information can be effectively reflected. From Fig. 11, we can observe that the MambaSCD architecture is not only capable of distinguishing changed features at different scales, but also accurately identifying the specific semantic categories of these changes.
@@ IV-D3 BDA
Table IX reports the accuracy of the MambaBDA architecture and some Transformer and CNN-based architectures on the xBD dataset. The performance of MambaBDA architecture significantly outperforms these representative BDA approaches. Compared to the current Transformer-based SOTA method, DamFormer [31], our three MambaBDA architectures have a 2.64, 4.12, and 4.39 higher , respectively. Moreover, it can be noticed that the accuracy improvement compared to DamFormer mainly comes from the more difficult damage classification task rather than the building localization task, which demonstrates that the proposed MambaBDA architecture can adequately learn the spatio-temporal relationships between the multi-temporal images so as to effectively distinguish different building damage levels.
In order to further show the misclassification between the different damage levels in more detail, we show their confusion matrices in Fig. 12. As can be seen from the confusion matrices, the categories “No damage” and “destroyed” are classified with high accuracy and are not easily misclassified into the remaining two categories. In contrast, the categories of “minor damage” and “major damage” are easily misclassified as other categories, especially the “minor damage” category. A large portion of the “minor damage” category is misclassified as “no damage” and “major damage”. There are mainly the following reasons. Firstly, the optical images in the xBD dataset were acquired with a relatively small viewing angle of the sensor. In many cases, “minor damage” is simply damage to the structure of the house. Little information can be obtained from the orthorectified optical image for such cases. In addition, there is a semantic ambiguity in the definition of the categories minor damage and major damage. We argue that it is difficult to address this problem by starting from the model architecture perspective. It can be considered to introduce of auxiliary information, such as height information or multi-view imagery [21, 69], etc. to help the model make more accurate classification.
@@
Table X lists the the number of parameters and GFLOPs of these BDA approaches. For MambaBDA-Small, our architecture yields more accurate building damage assessment results compared to Siamese-UNet, ChangeOS-101 and MTF, with an improvement in from 5.64 to 9.46 for a similar number of parameters and lower GFLOPs.
Fig. 13 shows the visualization results of building damages caused by different disaster events obtained by MambaBDA on the xBD dataset. MambaBDA can accurately localize and then differentiate the damage levels of buildings despite the varying type of building and the disaster event, implying the potential of the MambaBDA architecture for practical disaster response applications.
IV-E Different Spatio-Temporal Fusion Methods
@@
Modelling spatio-temporal relationships is significant to CD tasks [46, 17, 24]. To demonstrate the effectiveness of our proposed modelling mechanisms, we compare it with several commonly used spatial-temporal modeling methods in CD. They are concatenation operations with feature pyramid network (FPN) [24] (baseline), temporal-symmetric transformation based on 3D convolutional layers (TST) [24], RNN-based method [17], and Tranformer-based method [28]. Table XI lists the performance of these modeling methods on the three CD subtasks. On the simpler BCD tasks, Transformer does not yield significant advantages in performance compared to RNN. However, on the more difficult SCD and BDA tasks, the advantages of the Transformer architecture are illustrated. Compared with these current representative spatio-temporal relationship modeling approaches, our proposed modeling approach based on the Mamba architecture, which inherits the advantages of RNN and Transformer, shows superior performance on the three CD tasks of BCD, SCD, and BDA. Compared to the Transformer-based spatio-temporal modeling approach, our method improves the F1 value by 0.77 on BCD, SeK by 0.6 on the SCD task, and by 1.10 on the BDA task.
We further visualize the features extracted by the MambaBCD architecture in Fig. 14. In the features extracted by the encoder, the features in the region of high response values are not necessarily regions of interest, i.e., changed regions. Whereas, in the decoder, after the spatio-temporal relationships are gradually modeled by the proposed spatio-temporal modelling mechanisms, changed regions gradually exhibit high response values.
@@ IV-F Comparison to Other Backbone Networks
Finally, we simply compare the Mamba with some representative backbone networks in CD tasks on the SYSU dataset in Table XII. It can be seen that Mamba outperforms the two Transformer backbone networks, Swin-Transformer and MixFormer. This is because, in order to reduce the computational overhead of self-attention operation, these two Transformer backbone networks either adopt a local attention mechanism (Swin-Transformer) [34] or a strided convolutional layer (MixFormer) [35] to reduce the size of the feature maps, which will negatively affect the learning of global contextual information. The Mamba architecture, on the other hand, can reduce its computational consumption to a linear relationship with the number of tokens, thus eliminating the need for these methods, which cause information loss, and allowing the fully learning of global context information. As for the remote sensing change detection task, due to the multi-scale phenomenon of land-cover features and the differences in spatial resolution of different sensors, the sizes of changed objects often vary greatly. Therefore, the features extracted by Mamba through fully learning the global contextual information, are particularly important for the subsequent detection of these changed objects with different sizes.
@@ IV-G GPU Memory Footprint
In Section IV-D, we list the number of parameters and GFLOPs of different models. In this subsection, we further compare the GPU memory footprint of some models. Fig. 15 compares the GPU memory footprint of some CNN-based and Transformer-based methods with our MambaBCD architecture for varying input size. We can see that the trend of GPU memory footprint with input size for MambaBCD-Small is similar to some of the advanced CNN architectures (DSIFN and SNUNet). However, our MambaBCD is able to learn global spatial contextual information as well as adequately model spatio-temporal relationships, thereby exhibiting better detection performance. The GPU memory footprint of MambaBCD-Small is significantly lower compared to ChangeFormerV5, which employs the Transformer architecture as its backbone. Note that ChangeFormerV5 itself has employed strided convolutional layer to reduce the computational overhead of self-attention operations. Also, we can find that the lightweight architecture ChangeFormerV2 has surpassed the Mamba architecture and the two CNN architectures, DSIFN and SNUNet, in terms of the increasing trend of GPU memory footprint when the input size is increased from 1536 to 2048. These results imply that Transformer to be applicable to large scale scenarios of remote sensing images often comes at the cost of information loss. In comparison, Mamba architecture can avoid this problem.
@@ IV-H Robustness Against Degraded Input Data
Finally, we test the robustness of our method to degraded input data. In the inference phase, the data input to the model are subjected to different levels of Gaussian blurring, Gaussian noise, and scaling of the data by different proportions, respectively. Fig. 16 shows the degradation on F1 score of different methods with degraded data on the LEVIR-CD+ dataset. Firstly, we can see that the CNN-based methods are very poorly robust to Gaussian blurred data. F1 scores of SiamCRNN-101 and DSIFN drop by nearly 60 when increasing the standard deviation in the Gaussian blur from 0 to 2. BIT-101 also drops its F1 score by more than 34 due to the adoptation of ResNet to extract features. Whereas the two methods using Transformer as backbone, SwinTransUNet and ChangeFormerV5 are relatively robust to Gaussian blurred data, with F1 scores decreased by 14.83 and 18.43 , respectively. In contrast, the F1 score for our MambaBCD architecture drops by only 8.73.
@@
Fig. 16-(c) compares the robustness of models trained at a single scale to test data at different scales. The F1 value of MambaBCD decreases by only 2.88 when the scaling is increased from 1 to 1.5. When the scaling ratio decreases from 1 to 0.75, the F1 value of MambaBCD decreases by only 1.18. When the scaling factor is reduced to 0.5, all methods, including MambaBCD, show a larger decrease in F1 score. However, compared to the decline of 14.99 of SwinTransUNet and decline of 16.41 of ChangeFormerV5, the decline of MambaBCD decline is smaller, at 12.03 . All the above results again demonstrate the superiority of our proposed Mamba-based CD architecture.
@@ V Conclusion
In this paper, we explore the potential of the emerging Mamba architecture for remote sensing image change detection tasks. Three architectures, MambaBCD, MambaSCD, and MambaBDA are developed for binary change detection, semantic change detection, and building damage assessment tasks, respectively. The encoders of all three architectures are cutting-edge Visual Mamba architecture, which can adequately learn the global contextual information of the input image with linear complexity. Then, for the change decoder, in order to adequately learn spatio-temporal features, we propose three ways of modeling spatio-temporal relationships which can adequately exploiting the attributes and advantages of Mamba architecture. Experiments on five benchmark datasets have fully revealed the potential of Mamba for multi-temporal remote sensing image data processing tasks. Compared to the CNN-based and Transformer-based approaches, the proposed three architectures can achieve SOTA performance on their respective tasks. Compared to some commonly used backbone networks, the Mamba architecture can extract features that are more suitable for downstream CD tasks. The combination of our proposed three spatio-temporal relationship modeling mechanisms and the Mamba architecture provides a new perspective on modeling spatio-temporal relationships, showing much better detection results on all three subtasks. Additional experiments show that our Mamba-based architecture is robust to degraded input data.

### v3

@@ IV-B2 Evaluation Metrics
For evaluating model performance across different CD subtasks, we employ a set of metrics tailored to the specific requirements of BCD, SCD and BDA. In evaluating the BCD performance of the model, we use the following metrics:
@@
Cohen’s Kappa (KC): Assessing the agreement between two raters beyond chance, Cohen’s Kappa is given by
@@
For BDA, we follow the protocol of the xView2 Challenge [51] to split BDA into two subtasks, i.e., building localization and damage classification tasks. The following metrics are adopted to measure the performance of each subtask and the overall BDA performance:
F1 Score for Building Localization (): This metric assesses the model’s accuracy in the building localization task.
@@
Overall F1 Score (): This reflects the combined performance of the model in building location and damage classification tasks. It is defined as
@@ IV-C Comparison Methods
To show the superiority of our Mamba-based architecture in CD tasks, we select a number of state-of-the-art methods based on CNN and Transformer architectures to compare across three distinct tasks: BCD, SCD, and BDA.
For BCD, our comparison methods include a suite of CNN-based methods: FC-EF [22], FC-Siam-Diff [22], FC-Siam-Conc [22], SiamCRNN [17] (utilizing ResNet architectures ranging from ResNet-18 to ResNet-101 [67]), SNUNet [40], DSIFN [23], HANet [41], and CGNet [42]. For the Transformer-based methods, we compare against ChangeFormer (versions 1 to 6) [30], BIT (with ResNet-18 to ResNet-101 as the backbone for feature extraction) [28], TransUNetCD [33], and SwinSUNet [32].
@@ IV-D2 SCD
Based on the SCD results, the detailed “from-to” semantic change information can be obtained, which is presented in Fig. 10. Firstly, we can find that the “from-to” transition matrices obtained by our methods are very close to the ground truth, demonstrating the effectiveness of our MambaSCD architecture again. Then, based on the contents of the matrices, we can find that the three semantic changes with the highest percentage between the two temporal phases are “non-vegetated ground surface building”, “low vegetation building”, “low vegetation non-vegetated ground surface”. These results indicate that the areas covered by the SECOND dataset was in the process of urbanization during this period.
@@
Table X further lists the number of parameters and GFLOPs of these methods. Compared to the Transformer-based ScanNet method, the GFLOPs of our three architectures are lower. Compared to SMNet, another Transformer-based method, the proposed MamabaSCD-Tiny can achieve higher SeK values with a lower number of parameters and GFLOPs. These results demonstrate the effectiveness of the proposed methods and the potential of the Mamba architecture on the SCD task.
Fig. 11 shows some semantic change maps on the test set of the SECOND dataset. Due to land-cover category combinations, SCD often involves a considerable number of categories. For ease of visualization, we follow the visualization way in [24, 68, 20], that is, only plotting the land-cover categories of the changed regions in both pre- and post-event images. In this way, the “from-to” semantic change information can be effectively reflected. From Fig. 11, we can observe that the MambaSCD architecture is not only capable of distinguishing changed features at different scales but also accurately identifying the specific semantic categories of these changes.
@@ IV-D3 BDA
Table IX reports the accuracy of the MambaBDA architecture and some Transformer and CNN-based architectures on the xBD dataset. The performance of the MambaBDA architecture significantly outperforms these representative BDA approaches. Compared to the current Transformer-based SOTA method, DamFormer [31], our three MambaBDA architectures have a 2.64, 4.12, and 4.39 higher , respectively. Moreover, it can be noticed that the accuracy improvement compared to DamFormer mainly comes from the more difficult damage classification task rather than the building localization task, which demonstrates that the proposed MambaBDA architecture can adequately learn the spatio-temporal relationships between the multi-temporal images so as to effectively distinguish different building damage levels.
In order to further show the misclassification between the different damage levels in more detail, we show their confusion matrices in Fig. 12. As can be seen from the confusion matrices, the categories “no damage” and “destroyed” are classified with high accuracy and are not easily misclassified into the remaining two categories. On the contrary, the categories of “minor damage” and “major damage” are easily misclassified as other categories, especially the “minor damage” category. A large portion of the “minor damage” category is misclassified as “no damage” and “major damage”. There are mainly the following reasons. First, the optical images in the xBD dataset were acquired with a relatively small viewing angle of the sensor. In many cases, “minor damage” is nuance damages to the structure of the building. Little relevant information can be obtained from the nearly orthorectified optical image for such cases. Furthermore, there is a semantic ambiguity in the definition of the categories of minor damage and major damage. We argue that it is difficult to address this problem by starting from the model architecture perspective. It can be considered to introduce auxiliary information, such as height information or multi-view imagery [21, 69], etc. to help the model make a more accurate classification.
@@
Table X lists the number of parameters and GFLOPs of these BDA approaches. For MambaBDA-Small, our architecture yields more accurate results on building damage assessment compared to Siamese-UNet, ChangeOS-101, and MTF, with an improvement in from 5.64 to 9.46 for a similar number of parameters and lower GFLOPs.
Fig. 13 shows the visualization results of the building damage caused by different disaster events obtained by MambaBDA on the xBD dataset. MambaBDA can accurately localize and then differentiate the damage levels of buildings despite the varying type of building and the disaster event, implying the potential of the MambaBDA architecture for practical disaster response applications.
IV-E Different Spatio-Temporal Modeling Methods
@@
Modeling spatio-temporal relationships is significant for CD tasks [46, 17, 24]. To demonstrate the effectiveness of our proposed modeling mechanisms, we compare them with several commonly used spatial-temporal modeling methods in CD. They are concatenation operations with feature pyramid network (FPN) [24] (baseline), temporal-symmetric transformation based on 3D convolutional layers (TST) [24], RNN-based method [17], and Tranformer-based method [28]. Table XI lists the performance of these modeling methods on the three CD subtasks. On the simpler BCD tasks, Transformer does not yield significant advantages in performance compared to RNN. However, on the more difficult SCD and BDA tasks, the advantages of the Transformer architecture are illustrated. Compared with these current representative spatio-temporal relationship modeling approaches, our proposed modeling approach based on the Mamba architecture, which inherits the advantages of RNN and Transformer, shows superior performance on the three CD tasks of BCD, SCD, and BDA. Compared to the Transformer-based spatio-temporal modeling approach, our method improves the F1 value by 0.77 on BCD, SeK by 0.6 on the SCD task and by 1.10 on the BDA task.
We further visualize the features extracted by the MambaBCD architecture in Fig. 14. In the features extracted by the encoder, the features in the region of high response values are not necessarily regions of interest, i.e., changed regions. Whereas, in the decoder, after the spatio-temporal relationships are gradually modeled by the proposed spatio-temporal modeling mechanisms, changed regions gradually exhibit high response values.
@@ IV-F Comparison to Other Backbone Networks
Finally, we simply compare the Mamba with some representative backbone networks in CD tasks on the SYSU dataset in Table XII. It can be seen that Mamba outperforms the two Transformer backbone networks, Swin-Transformer and MixFormer. This is because, in order to reduce the computational overhead of self-attention operation, these two Transformer backbone networks either adopt a local attention mechanism (Swin-Transformer) [34] or a strided convolutional layer (MixFormer) [35] to reduce the size of the feature maps, which will negatively affect the learning of global contextual information. The adopted VMamba architecture, on the other hand, can reduce its computational consumption to a linear relationship with the number of tokens, thus eliminating the need for these methods, which cause information loss, and allowing for the fully learning of global context information. As for the remote sensing change detection task, due to the multi-scale phenomenon of land-cover features and the differences in spatial resolution of different sensors, the sizes of changed objects often vary greatly. Therefore, the features extracted by Mamba through fully learning the global contextual information are particularly important for the subsequent detection of these changed objects with different sizes.
@@ IV-G GPU Memory Footprint
In Section IV-D, we list the number of parameters and GFLOPs of different models. In this subsection, we further compare the GPU memory footprint of some models. Fig. 15 compares the GPU memory footprint of some CNN-based and Transformer-based methods with our MambaBCD architecture for varying input sizes. We can see that the trend of GPU memory footprint with input size for MambaBCD-Small is similar to some of the advanced CNN architectures (DSIFN and SNUNet). However, our MambaBCD is able to learn global spatial contextual information as well as adequately model spatio-temporal relationships, thereby exhibiting better detection performance. The GPU memory footprint of MambaBCD-Small is significantly lower compared to ChangeFormerV5, which employs the Transformer architecture as its backbone. Note that ChangeFormerV5 itself has employed strided convolutional layers to reduce the computational overhead of self-attention operations. Also, we can find that the lightweight architecture ChangeFormerV2 has surpassed the Mamba architecture and the two CNN architectures, DSIFN and SNUNet, in terms of the increasing trend of GPU memory footprint when the input size is increased from 1536 to 2048. These results imply that Transformer to be applicable to large-scale scenarios of remote sensing images often comes at the cost of information loss. In comparison, Mamba architecture can avoid this problem.
@@ IV-H Robustness Against Degraded Input Data
Finally, we test the robustness of our method to degraded input data. In the inference phase, the data input to the model is subjected to different levels of Gaussian blurring, Gaussian noise, and scaling of the data by different proportions, respectively. Fig. 16 shows the degradation of the F1 score of different methods with degraded data on the LEVIR-CD+ dataset. First, we can see that the CNN-based methods are very poorly robust to Gaussian blurred data. F1 scores of SiamCRNN-101 and DSIFN drop by nearly 60 when increasing the standard deviation of the Gaussian blur from 0 to 2. BIT-101 also drops its F1 score by more than 34 due to the adoption of ResNet to extract features. Whereas the two methods using Transformer as the backbone, SwinTransUNet and ChangeFormerV5 are relatively robust to Gaussian blurred data, with F1 scores decreasing by 14.83 and 18.43 , respectively. In contrast, the F1 score for our MambaBCD architecture drops by only 8.73.
@@
Fig. 16-(c) compares the robustness of models trained at a single scale to test data at different scales. The F1 value of MambaBCD decreases by only 2.88 when the scale increases from 1 to 1.5. When the scaling ratio decreases from 1 to 0.75, the F1 value of MambaBCD decreases only by 1.18. When the scaling factor is reduced to 0.5, all methods, including MambaBCD, show a larger decrease in the F1 score. However, compared to the decline of 14.99 of SwinTransUNet and the decline of 16.41 of ChangeFormerV5, the decline of the MambaBCD decline is smaller, at 12.03 . All the above results again demonstrate the superiority of our proposed Mamba-based CD architecture.
@@ V Conclusion
In this paper, we explore the potential of the emerging Mamba architecture for remote sensing image change detection tasks. Three architectures, MambaBCD, MambaSCD, and MambaBDA are developed for binary change detection, semantic change detection, and building damage assessment tasks, respectively. The encoders of all three architectures are cutting-edge VMamba architecture, which can adequately learn the global contextual information of the input image with linear complexity. Then, for the change decoder, in order to adequately learn spatio-temporal features, we propose three ways of modeling spatio-temporal relationships that can adequately exploiting the attributes and advantages of Mamba architecture. Experiments on five benchmark datasets have fully revealed the potential of Mamba for multi-temporal remote sensing image data processing tasks. Compared to the CNN-based and Transformer-based approaches, the proposed three architectures can achieve SOTA performance on their respective tasks. Compared to some commonly used backbone networks, the Mamba architecture can extract features that are more suitable for downstream CD tasks. The combination of our proposed three spatio-temporal relationship modeling mechanisms and the Mamba architecture provides a new perspective on modeling spatio-temporal relationships, showing much better detection results on all three subtasks. Additional experiments show that our Mamba-based architecture is robust to degraded input data.